Find the equations to the sides of triangles the coordinates of whose angular points are (1,4), (2,-3) and (-1,-2)

Grade:10

1 Answers

Aditya Gupta
4 years ago
i think you are unaware of the standard forms of eqns of a st line.
the two point form of a st line is given as:
y – y1= [(y2 –  y1)/(x2 – x1)]*(x – x1), this line represents the eqn of a line passing through P(x1, y1) and Q(x2, y2).
in above ques, we have to find the equations to the sides of triangles the coordinates of whose angular points are (1,4), (2,-3) and (-1,-2). 
lets first find equation to the side of triangle passing through (1,4) and (2,-3).
y – 4= [(-3 – 4)/(2 – 1)]*(x – 1)
or y+7x= 11
similarly you can find other sides eqns
kindly approve :=)